# Understanding Oversmoothing in Diffusion-Based GNNs From the Perspective of Operator Semigroup Theory

### v1

@@ Abstract
This paper presents a novel study of the oversmoothing issue in diffusion-based Graph Neural Networks (GNNs). Diverging from extant approaches grounded in random walk analysis or particle systems, we approach this problem through operator semigroup theory. This theoretical framework allows us to rigorously prove that oversmoothing is intrinsically linked to the ergodicity of the diffusion operator. This finding further poses a general and mild ergodicity-breaking condition, encompassing the various specific solutions previously offered, thereby presenting a more universal and theoretically grounded approach to mitigating oversmoothing in diffusion-based GNNs. Additionally, we offer a probabilistic interpretation of our theory, forging a link with prior works and broadening the theoretical horizon. Our experimental results reveal that this ergodicity-breaking term effectively mitigates oversmoothing measured by Dirichlet energy, and simultaneously enhances performance in node classification tasks.
Machine Learning, ICML
1 Introduction
@@
Graph Neural Networks (GNNs) have emerged as a powerful tool for learning graph-structured data, finding applications in various domains such as materials science (Merchant et al., 2023; Hestroffer et al., 2023), bioinformatics (Zhang et al., 2023b; Gao et al., 2023b; Strokach et al., 2020), and recommendation systems (Zhang et al., 2023a; Gao et al., 2023a). In recent years, continuous GNNs (Xhonneux et al., 2020) are proposed to generalize previous graph neural networks with discrete dynamics to the continuous domain by Neural ODE (Chen et al., 2018). Graph Diffusion Chamberlain et al. (2021b); Song et al. (2022); Choi et al. (2023) further extended the message-passing mechanism in classic GNNs under a partial differential equation (PDE) perspective. These works have made significant progress in terms of interpretability, stability, heterogeneous graphs, and beyond.
@@
Despite these advances, a fundamental challenge of GNNs lies in the phenomenon of oversmoothing (Oono & Suzuki, 2020), where repetitions of message passing may cause node representations to become indistinguishable and thus lose their discriminative power. For GNNs with discrete dynamics, several works (Xu et al., 2018; Zhao & Akoglu, 2020; Chen et al., 2020; Rusch et al., 2023) are proposed to relieve the oversmoothing issues. Additionally, recent works (Rusch et al., 2022; Thorpe et al., 2022; Wang et al., 2023) have verified the existence of oversmoothing in GNNs with continuous dynamics and most of them address this issue by introducing additional terms in graph diffusion equations, such as source term (Thorpe et al., 2022), Allen-Cahn term (Wang et al., 2023), and reaction term (Choi et al., 2023). However, such extra terms to graph diffusion are often under specific physical scenarios without a generic and unified overview, resulting in case-specific solutions with narrow applicability.
@@
In this paper, we propose a unified framework using operator semigroup theory to address this limitation. By viewing node features as solutions to the Cauchy problem associated with linear graph diffusion, we provide an in-depth understanding of the dynamics leading to oversmoothing. Building on this foundation, we propose a general and mild ergodicity-breaking condition, which accommodates specific solutions from previous research and further offers a more universal rule for designing terms to mitigate oversmoothing in diffusion-based GNNs.
@@
In summary, this paper makes several key contributions to the study of oversmoothing problem:
  • We introduce a comprehensive framework based on operator semigroup theory to analyze the oversmoothing issue in diffusion-based GNNs and provide a clear, theoretical pathway for addressing it.
  • Our work proposes an ergodicity-breaking condition that not only addresses the oversmoothing problem but also encompasses several specific extra terms identified in prior works, demonstrating its broad applicability.
  • We provide a probabilistic interpretation of our method, thereby establishing a connection with previous theoretical analyses and enriching the overall understanding of diffusion-based GNNs dynamics.
  • We substantiate our theoretical results through synthetic and real-world experiments.
1.1 Related Work
@@
Diffusion-based GNNs. Treating GNNs as the discretization of the continuous dynamical system is a rapidly growing sub-field of graph representation learning (Chamberlain et al., 2021b; Chen et al., 2022; Behmanesh et al., 2023; Wu et al., 2023). Since the message passing (MP) mechanism shows an intrinsic link to the diffusion process, several diffusion-based GNNs (Eliasof et al., 2021; Fu et al., 2022; Song et al., 2022) are proposed and conducted using ODE solver (Chen et al., 2018). GRAND (Chamberlain et al., 2021b) parameterizes the underlying graph diffusion process to learn the node representations. BLEND (Chamberlain et al., 2021a) considers graph as the discretization of a manifold and jointly conducts continuous feature learning and topology evolution based on the Beltrami flow.
@@
Oversmoothing. Oversmoothing refers to the effect that node features of graph neural networks (GNNs) tend to become more similar with the increase of the network depth, constraining the model expressive power for many GNNs (Thorpe et al., 2022; Oono & Suzuki, 2020). Many previous GNN models aim at overcoming oversmoothing (Xu et al., 2018; Chen et al., 2020; Zhao & Akoglu, 2020). GRAND++ (Thorpe et al., 2022) alleviates the oversmoothing problem by introducing the source term. Several GNNs quantitatively tackle the oversmoothing problem by analyzing Dirichlet energy (Rusch et al., 2022, 2023). Based on the Allen-Cahn particle system with repulsive force, ACMP (Wang et al., 2023) shows adaption for node classification tasks with high homophily difficulty. GREAD (Choi et al., 2023) introduces the reaction diffusion equation, encompassing different types of popular reaction equations, and empirically mitigating oversmoothing. Our method provides a more general framework and a more theoretically grounded method to address the oversmoothing problem.
1.2 Organization
In Sec.2, we provided a brief overview of graph diffusion and oversmoothing. In Sec.3, we introduced the operator semigroup and generator of graph diffusion, which serve as mathematical tools to describe the dynamics of diffusion-based GNNs. From the perspective of operator semigroups, we investigated oversmoothing in Sec.4. Theorem 4.3 establishes the connection between oversmoothing and the ergodicity of generator. Based on this result, Theorem 4.5 presents a method to address oversmoothing. In Sec.5, we complement the content of Sec.4 with an intuitive probabilistic perspective. Theorem 5.3 shows that the method proposed in Theorem 4.5 has a well-defined probabilistic interpretation.
2 Background
@@
Graph Neural Networks. GNNs have emerged as a powerful framework for learning representations of graph-structured data. MPNN (Gilmer et al., 2017) incorperates two stages: aggregating stage and updating stage. The core idea behind MP is the propagation of information across the graph. Inspired by MPNN, many GNN models typically operate by iteratively aggregating and updating node-level features based on the features of neighboring nodes. GCNs (Kipf & Welling, 2017) and GATs (Velickovic et al., 2017) are two prominent representatives in the field. GCNs leverage graph convolutional layers to capture and model relational information. GATs introduce the attention mechanism, allowing nodes to selectively attend to their neighbors during information propagation.
Graph Diffusion. Neural Ordinary Diffusion Equations (Neural ODEs) (Chen et al., 2018) present an innovative approach in constructing continuous-time neural network, where the derivative of the hidden state is parametrized using the neural networks. Chamberlain et al. (2021b) extends this framework to construct continuous-time Graph Neural Networks (GNNs). Graph diffusion is the process through which features propagate across the nodes of a graph over time. More formally, the graph diffusion on a graph $\mathcal{G}$ is defined as follows:
$\frac{\partial H(t)}{\partial t}:=[A(H(t))-I]H(t),$ (1)
$\frac{\partial h(u,t)}{\partial t}:=\sum_{v\in\mathcal{V}}a(h(u,t),h(v,t))(h(v,t)-h(u,t)),$ (2)
where $A(H(t))=(a(h(u,t),h(v,t)))_{u,v\in\mathcal{V}}$ is the attention matrix, satisfying the normalization condition: $\sum_{v\in\mathcal{V}}a(h(u,t),h(v,t))=1,\forall u$. When $A(H(t))$ solely relies on initial features, implying $A(H(t))=A$, Eq. (1) assumes the linear form:
$\frac{\partial h(u,t)}{\partial t}:=\sum_{v\in\mathcal{V}}a(u,v)(h(v,t)-h(u,t)).$ (3)
Moreover, by introducing a time discretization scheme, we can elucidate the forward propagation of the diffusion equation using the ordinary differential equation (ODE) solver. For instance, employing the Euler’s method with a time step of $\delta$, we arrive at:
$H(\delta(k+1))-H(\delta k)=\delta(A-I)H(\delta k)$
It is worth emphasizing that, many existing GNN models can be viewed as realization of the graph diffusion model by setting the feasible step and attention coefficients (Chamberlain et al., 2021b). This linkage allows us to leverage the properties of ODEs to analyze a diverse range of graph neural network models.
In GNNs, oversmoothing problem refers to the phenomenon where node features become excessively similar along with the propagation process. In the context of graph diffusion, the oversmoothing problem can be defined as follows:
Definition 2.1 (The oversmoothing problem ).
@@
3 Operator Semigroup and Generator of Graph Diffusion
In this section, we offer an extensive introduction to pertinent operator semigroup theory and associated theorems. These theoretical concepts are fundamental mathematical tools employed for the comprehensive analysis of the sustained behavior on graphs exhibited by dynamical systems over an extended period.
In the context where $\mathcal{V}=\{1,2,\ldots,n\}$, let $f:\mathcal{V}\rightarrow\mathbb{R}$ be some real-valued functions defined on node set $\mathcal{A}$ and $C(\mathcal{V})$ represent the set of all real-valued function on the node set. The operator is a mapping from function to function. The operator semigroup denoted as $\mathbf{S}=\{S_{t}\}_{t\geq 0}$ comprises a family of operators exhibiting the semigroup property: $S_{r}\circ S_{t}=S_{t}\circ S_{r}=S_{t+r}$, for $t,r\geq 0,$ and $S_{0}$ is an identity operator. Here $\circ$ denotes the composition operation.
Given operator semigroup $S_{t}$, operator
$\mathcal{L}=\lim\limits_{r\to 0}\frac{1}{r}\left[S_{r}-\mathcal{I}\right]$ (4)
is called the infinitesimal generator of $\mathbf{S}$, where $\mathcal{I}$ is the identity operator. $\mathcal{L}$ is uniquely determined given semigroup $S_{t}$The linearity of the operators $S_{t},t\geq 0$, together with the semigroup property, shows that $\mathcal{L}$ is the derivative of $S_{t}$ at any time $t>0$. for $t,r>0$,
$\frac{1}{r}\left[S_{t+r}-S_{t}\right]=S_{t}\left(\frac{1}{r}\left[S_{r}-\mathcal{I}\right]\right)=\left(\frac{1}{r}\left[S_{r}-\mathcal{I}\right]\right)S_{t}.$
Let $r\rightarrow 0$, yielding
$\frac{\partial S_{t}}{\partial t}=\mathcal{L}S_{t}=S_{t}\mathcal{L}.$ (5)
Eq. (5) is known as the Fokker-Planck equation.
Remark 3.1.
Conversely, given operator $\mathcal{L}$, the corresponding semigroup denoted as $S_{t}$ can also be uniquely determined from the Fokker-Planck equation.
@@ 3 Operator Semigroup and Generator of Graph Diffusion
$\mathcal{A}f(u):=\sum_{v\in\mathcal{V}}a(u,v)(f(v)-f(u)),$ (6)
Then operator semigroup of graph diffusion induced by $\mathcal{A}$ can be well defined using Fokker-Plank equation, denoted as $\mathbf{P}=\{P_{t},t\geq 0\}$. The uniqueness of $\mathbf{P}$ is ensured by Remark 3.1. If $f$ represents a $d$-dimensional function, the operator $\mathcal{A}$ remains well-defined. It can be perceived as operating independently on each dimension.
Leveraging the definition of graph diffusion, $\mathbf{P}$ can be shown to hold the following properties:
Proposition 3.2.
@@
See proofs in Appendix A.1. Property (i) is commonly known as the Markov property, meaning that the graph diffusion semigroup $P_{t}$ can be classified as a Markov semigroup. Furthermore, $\mathbf{P}$ has a precise probabilistic interpretation, as $P_{t}$ acts as a transition function for a Markov chain for every $t>0$. In the context of graph diffusion, $P_{t}$ represents an $n\times n$ transition probability matrix, where the row sums equal 1 and the transition probabilities between non-adjacent nodes are zero. For a more comprehensively probabilistic interpretation, please refer to Sec. 5.
4 Operator View for Oversmoothing
In this section, we illustrate that the graph diffusion equation (3), can be formulated as the Cauchy problem, and the solution of the graph diffusion equation has an operator semigroup format. We further discern that oversmoothing problem arises from the ergodicity of operators. To elaborate, the hidden layer features $h(u,t)$ converge to a uniform limit as $t\rightarrow\infty$, across all nodes $u\in\mathcal{V}$. To address this issue, we introduce a condition aimed at disrupting ergodicity and hence mitigating the oversmoothing problem, offering a more universal rule for designing terms to mitigate oversmoothing in diffusion-based GNNs.
4.1 Operator Representation of Feature
The dynamic of GRAND is given by the graph diffusion equation (Chamberlain et al., 2021b). Concretely, given the node input features $X(u)=[X_{1}(u),X_{2}(u),\ldots,X_{d}(u)]$, where $u\in\mathcal{V}$, the node features $h(u,t)$ are the solution to the following Cauchy problem
$\begin{cases}\frac{\partial h_{i}(u,t)}{\partial t}:=\mathcal{A}h_{i}(u,t),\\ h_{i}(u,0)=f_{i}(u).\end{cases}\quad\forall u\in\mathcal{V},$ (7)
where $f_{i}(u)=X_{i}(u)$ is the initial condition. The Fokker-Planck Equation (5) provides us with the solution to the Cauchy problem, leading to the following theorem:
Theorem 4.1.

### v2

@@ Abstract.
This paper presents an analytical study of the oversmoothing issue in diffusion-based Graph Neural Networks (GNNs). Generalizing beyond extant approaches grounded in random walk analysis or particle systems, we approach this problem through operator semigroup theory. This theoretical framework allows us to rigorously prove that oversmoothing is intrinsically linked to the ergodicity of the diffusion operator. Relying on semigroup method, we can quantitatively analyze the dynamic of graph diffusion and give a specific mathematical form of the smoothing feature by ergodicity and invariant measure of operator, which improves previous works only show existence of oversmoothing. This finding further poses a general and mild ergodicity-breaking condition, encompassing the various specific solutions previously offered, thereby presenting a more universal and theoretically grounded approach to relieve oversmoothing in diffusion-based GNNs. Additionally, we offer a probabilistic interpretation of our theory, forging a link with prior works and broadening the theoretical horizon. Our experimental results reveal that this ergodicity-breaking term effectively mitigates oversmoothing measured by Dirichlet energy, and simultaneously enhances performance in node classification tasks.
Graph neural networks, Oversmoothing, Operator Semigroup
journalyear: 2025copyright: rightsretainedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709324isbn: 979-8-4007-1245-6/25/08ccs: Mathematics of computing Functional analysisccs: Mathematics of computing Stochastic processesccs: Computing methodologies Machine learning
1. Introduction
@@
Graph Neural Networks (GNNs) have emerged as a powerful tool for learning graph-structured data, finding applications in various domains such as materials science (Merchant et al., 2023), bioinformatics (Zhang et al., 2023; Zhu et al., 2023), and recommendation systems (Qin et al., 2024). In recent years, continuous GNNs (Xhonneux et al., 2020) are proposed to generalize previous graph neural networks with discrete dynamics to the continuous domain by Neural ODE (Chen et al., 2018). Graph diffusion (Chamberlain et al., 2021b; Song et al., 2022; Choi et al., 2023) further extended the message-passing mechanism in classic GNNs under a partial differential equation (PDE) perspective. These works have made significant progress in terms of interpretability, stability, heterogeneous graphs, and beyond.
@@
Despite these advances, a fundamental challenge of GNNs lies in the phenomenon of oversmoothing (Oono and Suzuki, 2020; Wu et al., 2023; Chen et al., 2023), where repetitions of message passing may cause node representations to become indistinguishable and thus lose their discriminative power. For GNNs with discrete dynamics, several works (Xu et al., 2018; Zhao and Akoglu, 2020; Chen et al., 2020; Rusch et al., 2023) are proposed to relieve the oversmoothing issues. Additionally, recent works (Rusch et al., 2022; Thorpe et al., 2022; Wang et al., 2023) have verified the existence of oversmoothing in GNNs with continuous dynamics and most of them address this issue by introducing additional terms in graph diffusion equations, such as source term (Thorpe et al., 2022), Allen-Cahn term (Wang et al., 2023), and reaction term (Choi et al., 2023). However, such extra terms to graph diffusion are often under specific physical scenarios without a generic and unified overview, resulting in case-specific solutions with narrow applicability.
@@
In this paper, we propose a unified framework using operator semigroup theory to perform a principled analysis, then address this limitation. By viewing node features as solutions to the Cauchy problem associated with linear graph diffusion, we provide an in-depth understanding of the connection between ergodicity of operator and oversmoothing. In contrast to previous work studying oversmoothing through smooth measures such as Dirichlet energy (Maskey et al., 2023; Wu et al., 2023), with operator semigroups we can not only confirm the existence of oversmoothing, but also make it clear that the speed of oversmoothing is exponential and dependent on the spectral gap of the operator. Further, the operator semigroup approach gives a specific form for the fixed point, which improves the work that merely know that the fixed point exists. We also discuss the nonlinear case through the weak ergodicity of the nonlinear graph diffusion operator. Building on this foundation, we propose a general and mild ergodicity-breaking condition, which accommodates specific solutions from previous research and further offers a more universal rule for designing terms to mitigate oversmoothing in diffusion-based graph neural networks.
@@
In summary, this paper makes several key contributions to the study of oversmoothing problem: (1) We introduce a comprehensive framework based on operator semigroup theory to analyze the oversmoothing issue in diffusion-based graph neural networks. From this framework, specific mathematical form of the smoothing feature and rate of convergence can be given by the quantity associated with graph diffusion operator; (2) Our work proposes an ergodicity-breaking condition that not only addresses the oversmoothing problem but also encompasses several specific extra terms identified in prior works, demonstrating its broad applicability; (3) We provide a probabilistic interpretation of our method, thereby establishing a connection with previous theoretical analyses and enriching the overall understanding of diffusion-based graph neural networks dynamics; (4) We substantiate our theoretical results through synthetic and real-world experiments.
1.1. Related Work
@@
Diffusion-based graph neural networks. Treating GNNs as the discretization of the continuous dynamical system is a rapidly growing sub-field of graph representation learning (Chamberlain et al., 2021b; Behmanesh et al., 2023). Since the message passing (MP) mechanism shows an intrinsic link to the diffusion process, several diffusion-based graph neural networks (Eliasof et al., 2021; Fu et al., 2022; Song et al., 2022) are proposed and conducted using ODE solver (Chen et al., 2018). GRAND (Chamberlain et al., 2021b) parameterizes the underlying graph diffusion process to learn the node representations. BLEND (Chamberlain et al., 2021a) considers graph as the discretization of a manifold and jointly conducts continuous feature learning and topology evolution based on the Beltrami flow.
@@
Oversmoothing. Oversmoothing refers to the effect that node features of graph neural networks tend to become more similar with the increase of the network depth, constraining the model expressive power for many graph neural networks (Thorpe et al., 2022; Oono and Suzuki, 2020). Many previous graph neural networks aim at overcoming oversmoothing (Xu et al., 2018; Chen et al., 2020; Zhao and Akoglu, 2020). GRAND++ (Thorpe et al., 2022) relieves the oversmoothing problem by introducing the source term. Several GNNs quantitatively tackle the oversmoothing problem by analyzing Dirichlet energy (Rusch et al., 2022, 2023). Based on the Allen-Cahn particle system with repulsive force, ACMP (Wang et al., 2023) shows adaption for node classification tasks with high homophily difficulty. GREAD (Choi et al., 2023) introduces the reaction diffusion equation, encompassing different types of popular reaction equations, and empirically mitigating oversmoothing. FROND (Kang et al., 2024) interprets the feature updating process on graphs as a non-Markovian random walk which converges to the stationary distribution at algebraic rate thus relieving oversmoothing. Our method provides a more general framework and a more theoretically grounded method to address the oversmoothing problem.
Figure 1. Paper Organization
1.2. Organization
In Sec. 2, we provides a succinct introduction to the concepts of graph diffusion and oversmoothing. In Sec. 3, we delve into the operator semigroups and generators of graph diffusion, which are pivotal mathematical constructs for modeling the dynamics of diffusion-based GNNs. Sec. 4 explores oversmoothing from the viewpoint of operator semigroups. In particular, Theorem 3 in Sec. 4.2 elucidates the link between oversmoothing and the ergodic properties of the generator. Building on this foundation, Theorem 7 in Sec. 4.3 proposes a novel approach to mitigate oversmoothing. Sec. 5 enhances the discussions in Sec. 4 by providing a complementary probabilistic perspective as the connection and generalization on previous works. Theorem 3 validates that the strategy introduced in Theorem 7 possesses a coherent probabilistic interpretation. The organization of the main results of this work is outlined in Figure 1.
2. Background
@@
Graph Neural Networks. GNNs have emerged as a powerful framework for learning representations of graph-structured data. Message passing neural networks (MPNNs) (Gilmer et al., 2017) incorperates two stages: aggregating stage and updating stage. The core idea behind MP is the propagation of information across the graph. Inspired by MPNNs, many graph neural networks typically operate by iteratively aggregating and updating node-level features based on the features of neighboring nodes. GCNs (Kipf and Welling, 2017) and GATs (Velickovic et al., 2018) are two prominent representatives in the field. GCNs leverage graph convolutional layers to capture and model relational information. GATs introduce the attention mechanism, allowing nodes to selectively attend to their neighbors during information propagation.
Graph Diffusion. Neural Ordinary Diffusion Equations (Neural ODEs) (Chen et al., 2018) present an innovative approach in constructing continuous-time neural network, where the derivative of the hidden state is parametrized using the neural networks. Chamberlain et al. (2021b) extends this framework to construct continuous-time graph neural networks. Graph diffusion is the process through which features propagate across the nodes of a graph. More formally, the graph diffusion on a graph $\mathcal{G}$ is defined as follows:
(1) $\frac{\partial H(t)}{\partial t}:=[A(H(t))-I]H(t),$
(2) $\frac{\partial h(u,t)}{\partial t}:=\sum_{v\in\mathcal{V}}a(h(u,t),h(v,t))(h(v,t)-h(u,t)),$
where $I$ is identity matrix, $A(H(t))=(a(h(u,t),h(v,t)))_{u,v\in\mathcal{V}}$ is the attention matrix (Velickovic et al., 2018), where
(3) $a(h(u,t),h(v,t))=\frac{\exp{(b^{(t)}_{u,v}})}{\sum_{w\in\mathcal{N}(u)}\exp{(b^{(t)}_{u,w})}},\;(u,v)\in\mathcal{E},$
satisfying $\sum_{v\in\mathcal{V}}a(h(u,t),h(v,t))=1,\;\forall u$, where $\Phi:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is attention function to compute the weight between $u$ and $v$, $b^{(t)}_{u,v}:=\Phi(h(u,t),h(v,t))$. When $A(H(t))$ solely relies on initial features, implying $A(H(t))=A$, Eq. (2) is in the linear form:
(4) $\frac{\partial h(u,t)}{\partial t}:=\sum_{v\in\mathcal{V}}a(u,v)(h(v,t)-h(u,t)).$
In this paper, we mainly concerned linear diffusion-based graph neural networks. Further discussion on nonlinear case is in section 4.2. Moreover, by introducing a time discretization scheme, we can elucidate the forward propagation of the diffusion equation using the ODE solver. It is worth emphasizing that many existing graph neural networks can be viewed as realization of the graph diffusion model by setting the feasible step and attention coefficients (Chamberlain et al., 2021b). This linkage allows us to leverage the properties of ODEs to analyze a diverse range of graph neural network models.
In graph neural networks, oversmoothing problem refers to the phenomenon where node features become excessively similar along with the propagation process. In the context of graph diffusion, the oversmoothing problem can be defined as follows:
Definition 1 (The oversmoothing problem).
@@
We do not define oversmoothing here by a smoothness metric such as Dirichlet energy and emphasize exponential convergence because the necessity of exponential convergence as a condition for defining oversmoothing remains subject to debate. We opted not to adopt a definition that might spark controversy but instead expressed oversmoothing in mathematical forms based on the most fundamental intuitive understanding (Thorpe et al., 2022; Nguyen et al., 2023; Keriven, 2022). Indeed, exponential convergence can also be demonstrated from our proofs.
3. Operator Semigroup and Generator of Graph Diffusion
In this section, we offer an extensive introduction to pertinent operator semigroup theory and associated theorems. These theoretical concepts are fundamental mathematical tools for the comprehensive analysis of the long time behavior of dynamical systems on graphs.
For $\mathcal{V}=\{1,2,\ldots,n\}$, let $f:\mathcal{V}\rightarrow\mathbb{R}$ defined on node set and $C(\mathcal{V})$ represent the set of all real-valued function on the node set. The operator is a mapping from function to function. The operator semigroup denoted as $\mathbf{S}=\{S_{t}\}_{t\geq 0}$ comprises a family of operators with semigroup property: $S_{r}\circ S_{t}=S_{t}\circ S_{r}=S_{t+r}$, for $t,r\geq 0,$ and $S_{0}$ is an identity operator. Given operator semigroup $S_{t}$, operator
(5) $\mathcal{L}=\lim\limits_{r\to 0}\frac{1}{r}\left[S_{r}-\mathcal{I}\right]$
is called the infinitesimal generator of $\mathbf{S}$, where $\mathcal{I}$ is the identity operator. $\mathcal{L}$ is uniquely determined given semigroup $S_{t}$The semigroup property of the operators $S_{t},t\geq 0$ shows that for $t,r>0$, $\frac{1}{r}\left[S_{t+r}-S_{t}\right]=S_{t}\left(\frac{1}{r}\left[S_{r}-\mathcal{I}\right]\right)=\left(\frac{1}{r}\left[S_{r}-\mathcal{I}\right]\right)S_{t}.$ Let $r\rightarrow 0$,
(6) $\frac{\partial S_{t}}{\partial t}=\mathcal{L}S_{t}=S_{t}\mathcal{L}.$
If $S_{t}$ is a Markov semigroup, equation  (6) is referred to as Kolmogorov’s backward equation.
@@ 3. Operator Semigroup and Generator of Graph Diffusion
(7) $\mathcal{A}f(u):=\sum_{v\in\mathcal{V}}a(u,v)(f(v)-f(u)),$
Then operator semigroup of graph diffusion induced by $\mathcal{A}$ can be well defined using Kolmogorov’s equation, denoted as $\mathbf{P}=\{P_{t},t\geq 0\}$. If $f$ represents a $d$-dimensional function, the operator $\mathcal{A}$ remains well-defined. It can be perceived as operating independently on each dimension. From the definition of graph diffusion, $\mathbf{P}$ can be shown to hold the following properties:
Proposition 1.
@@
Property (i) is commonly known as the Markov property, meaning that the graph diffusion semigroup $P_{t}$ can be classified as a Markov semigroup. Furthermore, $\mathbf{P}$ has a clear probabilistic interpretation, as $P_{t}$ acts as a transition function for a Markov chain. In the context of graph diffusion, $P_{t}$ represents an $n\times n$ transition probability matrix, where the row sums equal 1 and the transition probabilities between non-adjacent nodes are zero. For a more comprehensively probabilistic interpretation, please refer to Sec. 5.
4. Operator View for Oversmoothing
In this section, we illustrate that the graph diffusion equation (4), can be formulated as the Cauchy problem, and the solution of the graph diffusion equation has an operator semigroup format. We further discern that oversmoothing problem arises from the ergodicity of operators. In specific, the hidden layer features $h(u,t)$ converge to a fixed point as $t\rightarrow\infty$, for all nodes $u\in\mathcal{V}$. To address this issue, we introduce a condition aimed at breaking ergodicity and hence relieving the oversmoothing problem, offering a more universal rule for designing terms to relieve oversmoothing in diffusion-based GNNs.
4.1. Operator Representation of Feature
The dynamic of GRAND (Chamberlain et al., 2021b) is given by the graph diffusion equation. Given node input features $X(u)=[X_{1}(u),X_{2}(u),\ldots,X_{d}(u)],u\in\mathcal{V}.$ The node features $h(u,t)$ are the solution to the following Cauchy problem
(8) $\begin{cases}\frac{\partial h_{i}(u,t)}{\partial t}:=\mathcal{A}h_{i}(u,t),\\ h_{i}(u,0)=f_{i}(u).\end{cases}\quad\forall u\in\mathcal{V},$
where $f_{i}(u)=X_{i}(u),i=1,\ldots,d$ are initial values. The Kolmogorov’s equation (6) provides us with the solution to the Cauchy problem, leading to the following theorem:
Theorem 1.